# Deliberate Planning of 3D Bin Packing on Packing Configuration Trees

### v3

@@ 4.5 Recursive Packing for Large Problems
[…]
@@
As illustrated in Table 6, recursive packing excels on large problem scales, achieving consistently the best performance. Notably, as the problem scale $\bar{N}$ increases, its performance continues to improve. Among all tested methods, only recursive packing and LSAH exhibit such improvements as the problem scale grows, with recursive packing significantly outperforming LSAH. We visualize the large-scale packing results of recursive packing in Appendix D. For the smaller scale $\bar{N}=200$, both PCT and PCT behave satisfactorily. However, as $\bar{N}$ increases, PCT quickly deteriorates, confirming training instability brought by long-sequence decision making and underscoring the necessity of problem decomposition. PCT performs consistently across different problem scales, verifying its generalization ability.
@@
The results are summarized in Table 7. Locally evaluating sub-bins without inter-bin comparison leads to sub-optimal performance. The minimum surface area alternative, while it evaluates globally, does not directly correlate with the objective of maximizing space utilization. These results emphasize the importance of effectively integrating local solutions. In contrast, our spatial ensemble method leverages inter-bin comparison to obtain global solution, resulting in significant performance advantage across different $\bar{N}$.
@@
Figure 9: Packing performance for varying sub-problem scale thresholds $\tau$, tested on problem scale $\bar{N}=500$. As $\tau$ nears $\bar{N}$, both algorithms degrade to direct transfer. Spatial ensemble benefits from problem decomposition, with smaller $\tau$ improving performance, while maximum return shows the opposite.
@@
We explore the impact of the sub-problem decomposition threshold $\tau$ on final performance, with experiments conducted on $\bar{N}=500$. As observed in Figure 9, increasing $\tau$, which makes sub-problem scale approach the original problem, causes both recursive packing and maximum return to degrade to the direct generalization performance of the pre-trained policy. The performance of recursive packing improves with finer decomposition, while the maximum return integration declines. This highlights the importance of solution integration choice.
@@ 4.6 ToP Results on BPP Variations
We evaluate the effectiveness of our unified planning framework, ToP, across different BPP variations. We compare ToP with state-of-the-art algorithms for each setting to validate its consistent superiority. We first conduct comparisons on packing forms with online properties with $u>0$. For lookahead packing with $s=1,p>0$, we compare ToP with CDRL (Zhao et al., 2022b). For buffering packing with $s>1,p=0$, we compare it with TAP-NET++ (Xu et al., 2023). For general packing, where $s\geq 1,p\geq 0$, we adopt the O3DBP method proposed by Kagerer et al. (2023). Since most baselines operate in discrete domains, this experiment is conducted on the discrete dataset. Performance comparisons on the continuous-domain ICRA stacking challenge datasets are provided in Appendix C.
Table 8: Performance comparison across different BPP variations with online properties.
@@
Table 9: Performance comparisons on offline 3D-BPP.
[…]
Uti. $\uparrow$ Var. $\downarrow$ Num. $\uparrow$ Gap $\downarrow$
Gurobi $50$ 00 00 $76.8\%$ 11.5 $29.9$ 19.3%
RCQL $50$ 00 00 $77.8\%$ $3.4$ $19.9$ $18.3\%$
Attend2Pack $50$ 00 00 $87.6\%$ $4.4$ $26.0$ $8.0\%$
TAP-NET++ $50$ 00 00 $89.2\%$ 1.2 $28.0$ $6.3\%$
ToP $50$ 00 00 95.2% 0.2 36.3 0.0%
@@ 4.7 Real-World Packing Robot
We develop a real-world packing robot in an industrial warehouse, as presented in Figure 11. The system adopts an ABB IRB 6700 robotic arm with a 210kg load capacity. Boxes (items) are delivered via a conveyor belt, which stops upon box detection by a photoelectric sensor. A key challenge for robotic packing is safely and efficiently placing boxes into constrained spaces where objects are positioned closely. Even minor robot-object collisions can destabilize the stack, leading to serious production failure. Moreover, unlike laboratory packing scenarios with protective container walls (Yang et al., 2021a; Xu et al., 2023), industrial packing often omits such protection for efficiency, making the constrained placement problem (Choset et al., 2005) more challenging. Solving it with motion planning (Görner et al., 2019) requires precise scene modeling and is time-consuming, violating practical demands.
@@
Instead of motion planning, we adopt a simple top-down placement manner with a flexible modular gripper which can actively adjust its shape to satisfy constrained placements. Each module is equipped with multiple suction cups for gripping flat items. Modules vary in size, allowing the gripper to adjust its own shape based on the target box’s size, as illustrated in Figure 12 (a). The shape adjustment principle is simplified to maximize box coverage without exceeding the box’s top boundary, providing sufficient gripping force while avoiding robot-object collisions during placement, as demonstrated in Figure 12 (b) and (c). This ensures both efficiency and safety in industrial scenarios.
The warehouse stocks 8000 Stock Keeping Units (SKUs), including large-sized boxes ranging up to 80 $\times$ 80 $\times$ 60cm, with a maximum weight of 30kg. The unprotected pallet measures 120 $\times$ 100cm, and the maximum stack height is 140cm. An RGB-D camera (PhoXi 3D Scanner XL) captures the top surface of incoming boxes. Item detection and segmentation are performed using Mask R-CNN (He et al., 2017), achieving a 99.95% recognition success rate for boxes on the conveyor. The modular gripper consists of 165 suction cups (39mm diameter), adjustable via a rodless pneumatic cylinder, with a maximum suction force of $260kg/m^{2}$. It automatically adjusts its configuration while approaching the target box, ensuring no delay in the packing cycle. The gripper has a 40cm lifting stroke, so surrounding boxes must not exceed 40cm above the top surface of the target placement. Placement candidates violating this constraint are discarded.
@@
We model real-world transportation uncertainties through physics-based verification. Ideally, if a real-world disturbance set $d$ causes stack $\mathbf{B}$ to collapse, a physical simulator $\mathcal{E}$, as $\mathbf{B}$’s digital twin, should predict this and reject unstable placements. However, real-world disturbances cannot be captured or recorded. To this, we randomly sample multiple disturbance sets to evaluate the stability for each placement, as illustrated in Figure 14 (a). Each disturbance set is a combination of 10 randomly generated translations and rotations, with translations in range $[15,20]$cm along the $x$ and $y$ axes and rotations in range $[-10^{\circ},10^{\circ}]$ for the $z$ axis. The simulated linear and angular velocities are set to 6m/s and $30^{\circ}$/s, respectively. If any disturbance causes the stack to collapse, we consider a placement $l$ unacceptable and turn to its alternatives (Figure 14 (b)). To meet real-time decision-making requirements of industrial packing, we leverage the Isaac Gym simulator (Makoviychuk et al., 2021), which feature batch parallelism and GPU-based simulation acceleration. For each placement, we sample $k_{d}$ disturbance sets, and evaluate $k_{l}$ such placements simultaneously, as exhibited in Figure 14 (c). We select the top $k_{l}$ placements based on probabilities output by policy $\pi_{\theta}$.
@@
The conveyor can hold up to 3 boxes for the robot to grip, modeling the general setting of packing with $1\leq s\leq 3$ and $0\leq p\leq 2$. We adopt ToP for solving different BPP variations out of the box. The task terminates when the stack reaches height limitation or no stable placement is available. AGVs or forklifts then transport the stack, and the packing restarts on a new pallet. During placement, the robot maintains a 1.5cm gap between each box. This is necessary due to the sensitivity of packing to operation accuracy.
Table 10: Real-world transportation stability (%). Test-time evaluation are conducted with quasi-static equilibrium (Zhao et al., 2022b) and our physics-based verification. Each is evaluated with 20 packing episodes.
@@
After box $n$ is placed, the RGB-D camera (HIKROBOT MV-DB1300A) on the pallet detects its position. If any deviation from the planned position is observed, the internal description of $n$ is updated accordingly to inform subsequent placement decisions. The average packing cycle is 10 seconds per box. Without container protection, 363 packing episodes are conducted in real-world industrial production, packing 6891 boxes with an average of 19 boxes per pallet and a 57.4% space utilization for relatively large-size boxes. Throughout the packing production, no robot-object collisions occurred and no stacks collapsed during transportation. We summarize how our physics-aware randomization improves transportation stability in Table 10, and take the disturbance number $k_{d}=8$ for test-time stability evaluation. A visual example of real-world packing is shown in Figure 16, with additional results provided in Appendix D. A video showcasing the dynamic packing process—featuring active gripper adjustment and stable transportation, is included in Supplementary Material.
@@ 5 Conclusions and Discussions
We formulate the online 3D-BPP as a novel hierarchical representation—packing configuration tree (PCT). PCT is a full-fledged description of the state and action space of bin packing which makes the DRL agent easy to train and well-performing. We extract state features from PCT using graph attention networks which encode the spatial relations of all space configuration nodes. The graph representation of PCT helps the agent with handling online 3D-BPP with complex practical constraints, while the finite leaf nodes prevent the action space from growing explosively. We further discover the potential of PCT as a tree-based planner to deliberately solve packing problems of industrial significance, including large-scale packing and different BPP variations.
Our method surpasses all other online packing competitors and is the first learning-based method that solves online 3D-BPP with continuous solution space. It performs well even on item sampling distributions that are different from the training one. We also give demonstrations to prove that PCT is versatile in terms of incorporating various practical constraints. The PCT-driven planning excels across large problem scales and different BPP variations, with performance improving as the problem scales and decision variables increase. Our real-world packing robot operates reliably on unprotected pallets, densely packing at 10 seconds per box with an average of 19 boxes and a 57.4% space utilization for relatively large-size boxes.
@@
We see several important opportunities for future research. First, we use the default physical simulation parameters (Makoviychuk et al., 2021) to evaluate the stability of real-world packing, where the Sim2Real gap can lead to incorrect solutions being either discarded or retained. From both academic and industrial perspectives, we are highly interested in estimating physical parameters from real-world packing processes to achieve system identification, support more accurate stability evaluation and Real2Sim2Real packing policy learning (Li et al., 2025). Additionally, this work models items as rigid bodies. Exploring ways to incorporate the deformability of items or containers (Bahety et al., 2023) to achieve tighter packing would be an intriguing direction. Last but not least, this work expands the industrial applicability of PCT through planning approaches, with its capability still bounded by the pre-trained base model. Developing PCT as a foundation model (Firoozi et al., 2025) trained on diverse packing data, with zero-shot generalization to new packing scenarios, is an exciting and promising avenue for future research.

### v4

@@ 4.5 Recursive Packing for Large Problems
We validate the effectiveness of recursive packing for large problem scales. To keep the validation focusing on problem decomposition and solution integration, we follow the standard packing setup (Martello et al., 2000) of setting 2setting\,2. The method is tested on continuous domains with bin dimensions set to $S^{d}=1$. The packing scale $\bar{N}$ is maintained by sampling the item size from a uniform distribution $\mathcal{U}(0,(8/\bar{N})^{\frac{1}{3}})$. Sub-problems are decomposed recursively with a threshold $\tau$, and our spatial ensemble method integrates local solutions produced by pre-trained PCT models $\pi_{\theta}$. The policy $\pi_{\theta}$ is exposed to multiple distributions of item sizes during training, specifically, $\mathcal{N}(0.3,0.1^{2})$, $\mathcal{N}(0.1,0.2^{2})$, and $\mathcal{N}(0.5,0.2^{2})$. The principle for determining $\tau$ is to ensure that the scale of decomposed sub-problems is regular and familiar to $\pi_{\theta}$. To this end, we generate 2000 random sequences from $\mathcal{U}(0.1,0.5)$ to evaluate $\pi_{\theta}$, and find the 95th percentile of packed item number to be 30.4. This indicates that 95% of the sequences contain no more than 30.4 items, and we set $\tau=30$.
The experiments are conducted on $\bar{N}\in\{200,500,1000\}$. To date, $\bar{N}=500$ is the largest scale for learning-based online packing. For each $\bar{N}$, 100 test sequences are randomly generated, and all methods are tested on the same data. The test results are summarized in Table 6. We compare with baselines which can operate in the continuous domain, including BR (Zhao et al., 2021), OnlineBPH (Ha et al., 2017), and LSAH (Hu et al., 2017). Additionally, we evaluate the performance of the PCT model trained on the corresponding scale $\bar{N}$, labeled as PCT, and the performance of transferring a pre-trained PCT model $\pi_{\theta}$ to large problem scales, labeled as PCT.
@@
As illustrated in Table 6, recursive packing excels on large problem scales, achieving consistently the best performance. Notably, as the problem scale $\bar{N}$ increases, its performance continues to improve. Among all tested methods, only recursive packing and LSAH exhibit such improvements as the problem scale grows, with recursive packing significantly outperforming LSAH. We visualize the large-scale packing results of recursive packing in Appendix F. For the smaller scale $\bar{N}=200$, both PCT and PCT behave satisfactorily. However, as $\bar{N}$ increases, PCT quickly deteriorates, confirming training instability brought by long-sequence decision making and underscoring the necessity of problem decomposition. A computational cost analysis in Appendix D also demonstrates that directly training PCT at larger scales incurs prohibitive overhead. PCT performs consistently across different problem scales, verifying its generalization ability.
@@
The results are summarized in Table 7. Locally evaluating sub-bins without inter-bin comparison leads to sub-optimal performance. The minimum surface area alternative, while it evaluates globally, does not directly correlate with the objective of maximizing space utilization. These results emphasize the importance of effectively integrating local solutions. In contrast, our spatial ensemble method leverages inter-bin comparison to obtain a global solution, resulting in a significant performance advantage across different $\bar{N}$.
@@
Figure 9: Packing performance for varying sub-problem scale thresholds $\tau$, tested on average problem scale $\bar{N}=500$. Spatial Ensumble performance peaks around $\tau=30$, aligning with the sub-problem scale frequently encountered by the base policy.
Table 8: Quantitative performance of recursive packing with Spatial Ensumble for solution integration across different $\tau$.
$\tau$ 5 10 15 20 25 30 35 40 45
Uti. $\uparrow$ 51.0% 71.0% 77.7% 79.3% 79.8% 79.9% 79.8% 79.7% 79.5%
50 100 150 200 250 300 350 400 450 500
79.3% 78.2% 76.5% 75.9% 75.5% 75.1% 74.8% 74.4% 74.1% 74.5%
@@
We explore the impact of the sub-problem decomposition threshold $\tau$ on final performance, with experiments conducted on $\bar{N}=500$. As observed in Figure 9, increasing $\tau$, which makes the sub-problem scale approach the original problem, causes both recursive packing and maximum return to degrade to the direct generalization performance of the pre-trained policy. The performance of recursive packing improves with finer decomposition, while the maximum return integration declines. This highlights the importance of the solution integration choice. The performance advantage persists as $\tau$ approaches 30. Further reducing $\tau$ causes the sub-problem scale to diverge from those encountered during $\pi_{\theta}$ training, leading to a gradual performance decline, as shown in Figure 9. We provide quantitative results in Table 9 to detail the relationship between $\tau$ and final performance.
@@ 4.6 ToP Results on BPP Variations
We evaluate the effectiveness of our unified planning framework, ToP, across different BPP variations. We compare ToP with state-of-the-art algorithms for each setting to validate its consistent superiority. We first conduct comparisons on packing forms with online properties with $u>0$. For lookahead packing with $s=1,p>0$, we compare ToP with CDRL (Zhao et al., 2022b). For buffering packing with $s>1,p=0$, we compare it with TAP-NET++ (Xu et al., 2023). For general packing, where $s\geq 1,p\geq 0$, we adopt the O3DBP method proposed by Kagerer et al. (2023). Since most baselines operate in discrete domains, this experiment is conducted on the discrete dataset. Performance comparisons on the continuous-domain ICRA stacking challenge datasets are provided in Appendix D.
Table 9: Performance comparison across different 3D-BPP variations with online properties.
Method Sel. Prev. Un. Setting 1Setting\,1 Setting 2Setting\,2 Setting 3Setting\,3
Uti. $\uparrow$ Var. $\downarrow$ Num. $\uparrow$ Gap $\downarrow$ Time $\downarrow$ Uti. $\uparrow$ Var. $\downarrow$ Num. $\uparrow$ Gap $\downarrow$ Time $\downarrow$ Uti. $\uparrow$ Var. $\downarrow$ Num. $\uparrow$ Gap $\downarrow$ Time $\downarrow$
CDRL 11 99 >0>0 $82.5\%$ $6.0$ $32.3$ $0.7\%$ 5.3 $75.0\%$ $1.3$ $29.0$ $17.4\%$ 3.1 $68.4\%$ $5.3$ $27.3$ $19.2\%$ 5.5
ToP 11 99 >0>0 83.1% 5.6 32.5 0.0% 1.8 90.8% 1.1 35.3 0.0% 1.1 84.7% 1.9 33.1 0.0% 1.8
O3DBP 55 55 >0>0 $53.1\%$ $7.5$ $26.4$ $39.0\%$ 3.7 $60.4\%$ $3.6$ $29.5$ $35.3\%$ 3.8 $53.4\%$ $7.4$ $26.5$ $39.5\%$ 3.8
ToP 55 55 >0>0 87.1% 1.2 34.2 0.0% 2.0 93.3% 0.9 36.3 0.0% 1.3 88.3% 1.3 34.4 0.0% 2.1
TAP-NET++ $10$ 0 >0>0 $66.8\%$ $3.6$ $20.6$ $24.8\%$ $\mathbf{6.2\times 10^{-2}}$ $77.0\%$ $2.1$ $26.6$ $18.9\%$ $\mathbf{4.6\times 10^{-2}}$ $68.2\%$ $3.0$ $21.2$ $23.5\%$ $\mathbf{6.3\times 10^{-2}}$
ToP $10$ 0 >0>0 88.8% 1.2 34.8 0.0% 2.1 95.0% 0.3 37.1 0.0% 1.5 89.1% 1.6 35.1 0.0% 2.1
The comparison results across different packing settings are summarized in Table 9. ToP consistently delivers the best space utilization. We visualize ToP performance with varying selectable item number $s$ and previewed item number $p$ in Figure 10, based on experiments conducted on setting 2setting\,2. The heatmap clearly reveals that, as decision variables increase, the packing performance also improves. Visualizations of ToP results across different BPP variations can be found in Appendix F.
Table 10: Performance comparisons on offline 3D-BPP.
Method Sel. Prev. Un. Setting 2Setting\,2
Uti. $\uparrow$ Var. $\downarrow$ Num. $\uparrow$ Gap $\downarrow$ Time $\downarrow$
Gurobi $50$ 0 0 $76.8\%$ 11.5 $29.9$ 19.3% 20.1
RCQL $50$ 0 0 $77.8\%$ $3.4$ $19.9$ $18.3\%$ $\mathbf{1.3\times 10^{-1}}$
@@
ToP $50$ 0 0 95.2% 0.2 36.3 0.0% 1.8
@@ 4.7 Real-World Packing Robot
We develop a real-world packing robot in an industrial warehouse, as presented in Figure 11. The system adopts an ABB IRB 6700 robotic arm with a 210kg load capacity. Boxes (items) are delivered via a conveyor belt, which stops upon box detection by a photoelectric sensor. A key challenge for robotic packing is safely and efficiently placing boxes into constrained spaces where objects are positioned closely. Even minor robot-object collisions can destabilize the stack, leading to serious production failure. Moreover, unlike laboratory packing scenarios with protective container walls (Yang et al., 2021b; Xu et al., 2023), industrial packing often omits such protection for efficiency, making the constrained placement problem (Choset et al., 2005) more challenging. Solving it with motion planning (Görner et al., 2019) requires precise scene modeling and is time-consuming, violating practical demands.
@@
Instead of motion planning, we adopt a simple top-down placement manner with a flexible modular gripper which can actively adjust its shape to satisfy constrained placements. Each module is equipped with multiple suction cups for gripping flat items. Modules vary in size, allowing the gripper to adjust its own shape based on the target box’s size, as illustrated in Figure 12 (a). The shape adjustment principle is simplified to maximize box coverage without exceeding the box’s top boundary, providing sufficient gripping force while avoiding robot-object collisions during placement, as demonstrated in Figure 12 (b) and (c). This ensures both efficiency and safety in industrial scenarios. We provide more details of our modular gripper and its simplified motion planning in Appendix E.
The warehouse stocks 8000 Stock Keeping Units (SKUs), including large-sized boxes ranging up to 80 $\times$ 80 $\times$ 60cm, with a maximum weight of 30kg. The unprotected pallet measures 120 $\times$ 100cm, and the maximum stack height is 140cm. An RGB-D camera (PhoXi 3D Scanner XL) captures the top surface of incoming boxes. Box detection and segmentation are performed using Mask R-CNN (He et al., 2017), achieving a 99.95% recognition success rate for boxes on the conveyor. An eye-to-hand calibration (Dekel et al., 2020) is performed between the on-conveyor camera and the robot base to transform the top-center coordinates of the detected items into the robot’s coordinate system for grasping. The modular gripper automatically adjusts its configuration while approaching the target box, ensuring no delay in the packing cycle. The gripper has a 40cm lifting stroke, so surrounding boxes must not exceed 40cm above the top surface of the target placement. Placement candidates violating this constraint are discarded.
@@
We model real-world transportation uncertainties through physics-based verification. Ideally, if a real-world disturbance set $d$ causes stack $\mathbf{B}$ to collapse, a physical simulator $\mathcal{E}$, as $\mathbf{B}$’s digital twin, should predict this and reject unstable placements. However, real-world disturbances cannot be captured or recorded. To this, we randomly sample multiple disturbance sets to evaluate the stability for each placement, as illustrated in Figure 14 (a). Each disturbance set is a combination of 10 randomly generated translations and rotations, with translations in the range [15,20][15,20]cm along the $x$ and $y$ axes and rotations in the range $10^{\circ},10^{\circ}]$ for the $z$ axis. The simulated linear and angular velocities are set to 6m/s and $30^{\circ}$/s, respectively. If any disturbance causes the stack to collapse, we consider a placement $l$ unacceptable and turn to its alternatives (Figure 14 (b)). To meet real-time decision-making requirements of industrial packing, we leverage the Isaac Gym simulator (Makoviychuk et al., 2021), which features batch parallelism and GPU-based simulation acceleration. For each placement, we sample $k_{d}$ disturbance sets, and evaluate $k_{l}$ such placements simultaneously, as exhibited in Figure 14 (c). We select the top $k_{l}$ placements based on probabilities output by policy $\pi_{\theta}$.
@@
The conveyor can hold up to 3 boxes for the robot to grip, modeling the general setting of packing with $1\leq s\leq 3$ and $0\leq p\leq 2$. We adopt ToP for solving different BPP variations with no additional adaptation. The task terminates when the stack reaches the height limitation or no stable placement is available. AGVs or forklifts then transport the stack, and the packing restarts on a new pallet. During placement, the robot maintains a 1.5cm gap between each box. This is necessary due to the sensitivity of packing to operation accuracy.
Table 11: Real-world transportation stability (%). Test-side evaluation is conducted with quasi-static equilibrium (Zhao et al., 2022b) and our physics-based verification. Each is evaluated with 20 packing episodes.
@@
After box $n$ is placed, the RGB-D camera (HIKROBOT MV-DB1300A) on the pallet detects its position. If any deviation from the planned position is observed, the internal description of $n$ is updated accordingly to inform subsequent placement decisions. The average packing cycle is 9.8 seconds per box, with a variance of 0.17, indicating consistent performance across repeated trials. The minor variations in cycle time result from differing object placements, which lead to varying transportation distances during execution. Without container protection, 363 packing episodes are conducted in real-world industrial production, packing 6891 boxes with an average of 19 boxes per pallet and a 57.4% space utilization for large-size boxes. Throughout the packing production, no robot-object collisions occurred, and no stacks collapsed during transportation. We summarize how our physics-aware randomization improves transportation stability in Table 11, and take the disturbance number $k_{d}=8$ for test-time stability evaluation. A visual example of real-world packing is shown in Figure 16, with additional results provided in Appendix F. A video showcasing the dynamic packing process, featuring active gripper adjustment and stable transportation, is included in Supplementary Material.
@@ 5 Conclusions and Discussions
We formulate the online 3D-BPP as a novel hierarchical representation—packing configuration tree (PCT). PCT is a full-fledged description of the state and action space of bin packing, which makes the DRL agent easy to train and well-performing. We extract state features from PCT using graph attention networks which encode the spatial relations of all space configuration nodes. The graph representation of PCT helps the agent with handling online 3D-BPP with complex practical constraints, while the finite leaf nodes prevent the action space from growing explosively. We further discover the potential of PCT as a tree-based planner to deliberately solve packing problems of industrial significance, including large-scale packing and different BPP variations.
Our method surpasses all other online packing competitors and is the first learning-based method that solves online 3D-BPP with continuous solution space. It performs well even on item sampling distributions that are different from the training one. We also give demonstrations to prove that PCT is versatile in terms of incorporating various practical constraints. The PCT-driven planning excels across large problem scales and different BPP variations, with performance improving as the problem scales and decision variables increase. Our real-world packing robot operates reliably on unprotected pallets, densely packing at 10 seconds per box with an average of 19 boxes and a 57.4% space utilization for large-size boxes.
@@
We see several important opportunities for future research. First, we use the default physical simulation parameters (Makoviychuk et al., 2021) to evaluate the stability of real-world packing, where the Sim2Real gap can lead to incorrect solutions being either discarded or retained. From both academic and industrial perspectives, we are highly interested in estimating physical parameters from real-world packing processes to achieve system identification (Ljung, 1998), support more accurate stability evaluation and Real2Sim2Real packing policy learning (Li et al., 2025). Additionally, this work proves from a geometric perspective that the PCT candidates cover placements that are locally optimal in a tightness measure; however, it does not provide tight performance bounds for the parameterized policies. Previous study (Yang et al., 2021a) provides only a loose bound on the competitive ratio for the online 3D-BPP, showcasing that the performance gap between the offline optimum and any online methods is no smaller than $\Omega(\log\theta\alpha)$, where $\theta$ is the ratio between the maximum and minimum item value, and $\alpha$ is the ratio of the aggregate capacity to the minimum capacity. This bound primarily depends on the distribution of the input problem data, while tighter bounds are algorithm-specific. Establishing tight performance guarantees for parameterized PCT policies whose behavior evolves throughout training remains a challenging yet meaningful future direction, as it helps clarify the theoretical positioning of PCT. Last but not least, this work expands the industrial applicability of PCT through planning, with its capability still bounded by the pre-trained base model. Developing PCT as a foundation model (Firoozi et al., 2025) trained on diverse packing data, with zero-shot generalization to new packing scenarios, is an exciting and promising avenue for future research.
6 Acknowledgments
This work was supported in part by the NSFC (62325211, 62132021, 62502347, 62522219, 62372457, 62322207, 62225113), the Postdoctoral Fellowship Program and China Postdoctoral Science Foundation (BX20250386), the Major Program of Xiangjiang Laboratory (23XJ01009), Key R&D Program of Wuhan (2024060702030143).